# HunyuanVideo: A Systematic Framework For Large Video Generative Models

### v4

@@ 7.2.1 Pre-training
Figure 19: Differences between text-to-video (T2V) model and image-to-video (I2V) model.
Image-to-video (I2V) task is a common application in video generation tasks. It usually means that given an image and a caption, the model uses this image as the first frame to generate a video that matches the caption. Although the naive HunyuanVideo is a text-to-video (T2V) model, it can be easily extended to an I2V model. Specifically, as mentioned in Sec. 4.2, the T2V model’s input is a latent $X$ with a shape of $T\times C\times H\times W$, where $T$, $C$, $H$ and $W$ represent the frame, channel, height and width of the compressed video respectively. Similar to Emu [25], in order to introduce image condition $I$, we treat $I$ as the first frame of a video and apply zero-padding to create a $T\times C\times H\times W$ tensor $I_{o}$, as shown in Fig. 19. Additionally, we employ a binary mask $m$ with dimensions $T\times 1\times H\times W$, where the first temporal position is set to 1, and all other positions are set to zero. Then the latent $X$, the tensor $I_{o}$ and the mask $m$ are concatenated along the channel dimension to form the input for the model. Note that since the channel of the input tensor has changed from $C$ to $12C+$ 1, as shown in Fig. 19, we need to adjust the parameters of the first convolutional module of the model from $\phi=(C_{in}(=C),C_{out},s_{h},s_{w})$ to $\phi^{\prime}=(C_{in}^{\prime}(=2C+1),C_{out},s_{h},s_{w})$, where each component corresponds to the input channel $C_{in}/C_{in}^{\prime}$, output channel $C_{out}$, height of the convolutional kernel $s_{h}$, and width of the convolutional kernel $s_{w}$. In order to retain the representation ability of the T2V model, the first $C$ input channels of $\phi^{\prime}$ are directly copied to $\phi$, and the rest are initialized to zero. We pre-train the I2V model on the same data as T2V model, and the results are shown in Fig. 20.
@@ 8 Related Works
  • Contributors: Jihong Zhang, Meng Chen, Jianchen Zhu, Winston Hu, Yongming Rao, Kai Liu, Lifei Xu, Sihuan Lin, Yifu Sun, Shirui Huang, Lin Niu, Shisheng Huang, Yongjun Deng, Kaibo Cao, Xuan Yang, Hao Zhang, Jiaxin Lin, Chao Zhang, Fei You, Yuanbin Chen, Yuhui Hu, Liang Dong Zheng Fang, Dian Jiao, Zhijiang Xu

### v5

@@ 7.2.1 Pre-training
Figure 19: HunyuanVideo-I2V Diffusion Backbone.
Figure 20: The noisy input differences between text-to-video model and image-to-video model.
Image-to-video (I2V) task is a common application in video generation tasks. It usually means that given an image and a caption, the model uses this image as the first frame to generate a video that matches the caption. Although the naïve HunyuanVideo is a text-to-video (T2V) model, it can be easily extended to an I2V model, as shown in Fig. 19. To enhance the model’s capability to comprehend the semantics of the input image and to more effectively integrate the information from both the image and the caption, the I2V model incorporates a semantic image injection module. It first inputs the image into the MLLM model to obtain the semantic image token and then concatenates these tokens into the video latent token for full-attention calculation. Similar to Emu [25], the I2V model also utilizes the image latent concatenation method to assist the model in more accurately reconstructing the original image information in its output. Specifically, as mentioned in Sec. 4.2, the T2V model’s input is a latent $X$ with a shape of $T\times C\times H\times W$, where $T$, $C$, $H$ and $W$ represent the frame, channel, height and width of the compressed video respectively. In the I2V model, we treat $I$ as the first frame of a video and apply zero-padding to create a $T\times C\times H\times W$ tensor $I_{o}$, as shown in Fig. 20. Additionally, we employ a binary mask $m$ with dimensions $T\times 1\times H\times W$, where the first temporal position is set to 1, and all other positions are set to zero. Then the latent $X$, the tensor $I_{o}$, and the mask $m$ are concatenated along the channel dimension to form the input for the model. Note that since the channel of the input tensor has changed from $C$ to $12C+$ 1, as shown in Fig. 20, we need to adjust the parameters of the first convolutional module of the model from $\phi=(C_{in}(=C),C_{out},s_{h},s_{w})$ to $\phi^{\prime}=(C_{in}^{\prime}(=2C+1),C_{out},s_{h},s_{w})$, where each component corresponds to the input channel $C_{in}/C_{in}^{\prime}$, output channel $C_{out}$, height of the convolutional kernel $s_{h}$, and width of the convolutional kernel $s_{w}$. In order to retain the representation ability of the T2V model, the first $C$ input channels of $\phi^{\prime}$ are directly copied to $\phi$, and the rest are initialized to zero. We pre-train the I2V model on the same data as the T2V model, and the results are shown in Fig. 21.
@@ 8 Related Works
  • Contributors: Jihong Zhang, Meng Chen, Jianchen Zhu, Winston Hu, Yongming Rao, Kai Liu, Lifei Xu, Sihuan Lin, Yifu Sun, Shirui Huang, Lin Niu, Shisheng Huang, Yongjun Deng, Kaibo Cao, Xuan Yang, Hao Zhang, Jiaxin Lin, Chao Zhang, Fei You, Yuanbin Chen, Yuhui Hu, Liang Dong Zheng Fang, Dian Jiao, Zhijiang Xu, Xuhua Ren, Bing Ma, Jiaxiang Cheng, Wenyue Li, Tianxiang Zheng